# Gaussian Mapping for Evolving Scenes

### v1

@@ 4.3 Keyframe Management
Since jointly optimizing Gaussians using all frames from a video stream is computationally infeasible, we maintain a small window $W_{k}$ of keyframes based on inter-frame covisibility. Effective keyframe management aims to select non-redundant keyframes that observe the same region while spanning a wide baseline to enforce stronger multiview constraints. In GaME, keyframes are selected every time a frame exceeds a translation $\theta_{\text{translation}}\in\mathbb{R}$ or a rotation $\theta_{\text{rotation}}\in\mathbb{R}$ threshold.
@@
Covisibility. Upon triggering Add or Remove operations by DSA, all covisible keyframes are retrieved by reprojecting 3D points from the current frame and selecting frames where these points are not occluded. This approach differs from conventional 3DGS methods [18, 33], which determine covisibility based on the number of Gaussians used to re-render the keyframes. We found projection to be more reliable than re-rendering in multi-room environments, as Gaussians from different rooms may still appear visible due to the alpha blending, leading to erroneous covisibility estimates.
@@
Keyframe Masking. Importantly, in evolving scenes, excluding outdated visual information is crucial, as the scene may change over time, making previous observations detrimental for the 3DGS optimization process. However, discarding entire keyframes that observe stale geometry can result in losing useful information. While some parts of the scene may change, other regions often remain stable and can provide valuable signals for 3DGS optimization. Therefore, rather than discarding entire keyframes as stale, GaME selectively ignores only the stale areas within them. To achieve this, before removing $\mathcal{G}_{\text{remove}}$ or $\mathcal{G}_{\text{remove}}^{\text{KF}}$, they are rendered onto the keyframe, and the image regions rendering them are masked out. The same procedure is applied for $\mathcal{G}_{\text{add}}$. Formally, the masked region $\mathcal{M}^{\text{KF}}$ within a keyframe is defined as:
$\mathcal{M}^{\text{KF}}=\left\{\mathbf{p}\in\mathbb{Z}^{2}\mid\mathbf{p}\in\text{render}(\mathcal{G}_{\text{remove}})\cup\text{render}(\mathcal{G}_{\text{remove}}^{\text{KF}})\cup\text{render}(\mathcal{G}_{\text{add}})\right\}$ (13)
where $\mathbf{p}$ represents a pixel location within the keyframe, and $*\text{render}(\mathcal{G}_{\text{*}})$ denotes the set of pixels influenced by the Gaussian set $*\mathcal{G}_{\text{*}}$ when rendered. Keyframes dominated by stale regions are of little value for optimization. We exclude any keyframe in which the instance masks of objects are more than $\theta_{\text{ignore}}$ percent covered by inconsistent observations. During optimization, for all the keyframes that have ignored areas, GaME optimizes the losses:
$L_{\text{color}}(\hat{I},I)=\frac{1}{\sum_{p}\mathbb{I}[p\notin\mathcal{M}^{\text{KF}}]}\sum_{p}\mathbb{I}[p\notin\mathcal{M}^{\text{KF}}]\cdot\lVert\hat{I}(p)-I(p)\rVert,$ (14)
$L_{\text{depth}}(\hat{D},D)=\frac{1}{\sum_{p}\mathbb{I}[p\notin\mathcal{M}^{\text{KF}}]}\sum_{p}\mathbb{I}[p\notin\mathcal{M}^{\text{KF}}]\cdot|\hat{D}(p)-D(p)|,$ (15)
Note that compared to (5), the structure similarity term is dropped as its convolutional nature does not interact well with masking.
Refinement. An added benefit of GaME is that, optionally, the masked keyframes can be directly used for further refinement after the mapping operation, leading to consistent optimization despite changes occurring throughout the mapping stage.
@@
SplaTAM [11] MonoGS [18] DG-SLAM [33] GaME (Ours) Ground Truth
@@ 5 Experiments
Datasets. We test our method on the Flat [26] dataset, which consists of two RGB-D sequences captured in a synthetic environment with significant changes occurring between. We further evaluate on the Aria [20] dataset to assess performance on real-world data. We select two recordings from two rooms each that have undergone long-term changes. Finally, the TUM-RGBD [28] dataset is used to assess performance on three different static scenes following the protocol of [18].
Evaluation Metrics. To assess rendering quality, we compute PSNR, SSIM [30] and LPIPS [38]. Rendering metrics on all the datasets is evaluated by rendering full-resolution images along the ground-truth trajectory. We assess the depth error using the L1 norm in centimeters.
Baselines. We compare GaME with state-of-the-art 3DGS online reconstruction systems MonoGS [18] and SplaTAM [11]. We compare to DG-SLAM [16], a recent dynamic 3DGS method, to assess the ability of dynamic 3DGS SLAM to handle evolving scenes.
@@ 5.1 Evolving Scene Mapping Results
We run GaME on the Flat and Aria datasets to evaluate rendering quality in evolving scenes, shown in Tabs. 1 and 2. The Flat dataset is designed to test mapping under scene changes and includes more substantial long-term dynamics. We note that all losses are computed over complete images, where changed areas typically occupy a smaller part. Nonetheless, we observe notable differences in rendering performance on both datasets, where GaME is the only method that accurately adapts the reconstruction to even subtle scene changes without compromising rendering quality. This granularity is shown in qualitative results in Fig. 3. Beyond high rendering quality, GaME is able to accurately resolve changes also for challenging cases such as small cutlery on the table and flat paintings.
@@
Table 1: Rendering performance on the Flat dataset. GaME shows superior performance in both color and depth rendering. Cells show metrics for input / novel views.
@@
DG-SLAM [16] room0 15.78 / 15.63 0.578 / 0.580 0.761 / 0.763 67.60 / 69.05
room1 12.62 / 12.44 0.651 / 0.643 0.702 / 0.708 55.47 / 57.03
Avg 14.20 / 14.04 0.615 / 0.612 0.732 / 0.736 61.54 / 63.04
@@
Table 2: Rendering performance on the Aria dataset. GaME shows superior performance in both color and depth rendering. Cells show metrics for input / novel views.
@@ 5.2 Ablation Studies
We conduct ablation studies to assess key design choices and analyze the reconstruction performance of both evolving and static scenes. Additional experiments on runtime and robustness to noisy input poses are provided in the Supplementary Material.
Add and Remove operations are important for Dynamic Scene Adaptation. We ablate the components of DSA in Tab. 3, showing that its presence significantly improves performance. While differences between individual DSA variants are less pronounced in final rendering quality, due to keyframe optimization correcting finer detail, combining both operations consistently yields the best performance and fastest adaptation.
[…]
Table 3: Dynamic Scene Adaptation (DSA) ablation in the Flat dataset. Combination of Add and Remove operations give the best performance. Cells show metrics for input / novel views.
@@
[…]
Table 4: Keyframe management ablation on Flat dataset. Retaining keyframes is essential to constrain the background. However, to avoid artifacts, conflicting regions must be accurately filtered.
GaME can reconstruct static scenes. While our method is specifically designed to reconstruct evolving scenes, it should not lose the ability to reconstruct static scenes. Results on the TUM-RGBD dataset shown in Tab. 5 show that GaME performs on par with state-of-the-art systems. This suggests that the introduced DSA and keyframe management are robust to noise and false positives in model updates and masking, which could deteriorate performance.
@@
Table 5: Rendering performance on TUM-RGBD dataset. GaME performs on par with state-of-the art mapping systems for static scenes. PSNR [dB]$\downarrow$ shown for input views.
GaME is an online mapping system. Our approach is of an incremental nature with partial optimization performed on every new keyframe. Optionally, our approach lends itself to further final refinement. Both options are shown in Tab. 6. While refinement can potentially improve the final rendering quality, we note that the model is already well converged after incremental processing.
Method PSNR [dB] $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Depth L1 [cm] $\downarrow$ No Refinement 24.48 / 23.30 0.899 / 0.883 0.231 / 0.266 17.00 / 15.90 With Refinement 24.76 / 23.47 0.898 / 0.881 0.229 / 0.267 16.90 / 15.80
Table 6: Final refinement ablation on Flat dataset. While refinement can potentially improve the final rendering quality, GaME already converges well during incremental processing.
@@ 6 Limitations
While GaME currently demonstrates robust rendering performance on evolving long-term dynamic scenes, several notable limitations exist. First, it does not yet handle short-term dynamic objects, which would result in inefficient addition and removal. Extending the dynamic scene adaptation mechanism to consider both dynamics is an exciting future direction. Second, we primarily study mapping with external pose tracking. Although sparse odometry systems can reject many changes as outliers [2], change-aware pose refinement and integration of a dedicated tracking mechanism is an exciting future direction that could improve accuracy. Third, while GaME is an online algorithm that can process sequences captured in evolving scenes, the current implementation has not been optimized for performance, which limits real-time deployment. Consistently updating the 3DGS map through incremental segmentation, rather than iterating over segmentation masks, could significantly speed up the approach. Finally, DSA is currently performed per incoming keyframe. This has the advantage of not storing semantic information in the 3DGS model and comes with some inherent robustness through the multi-view optimization process. However, individual semantic masks can be noisy or missing, where a delayed decision or global optimization could improve performance by considering more semantic observations.
@@ 7 Conclusions
We presented GaME, the first online mapping system for evolving scenes with novel view synthesis capabilities. GaME utilizes novel dynamic scene adaptation operations to detect and correct conflicts in the incrementally built 3DGS model. Our keyframe management method furthermore appropriately ignores stale areas in keyframes, retaining useful information while correcting for changed regions, resulting in a well-conditioned 3DGS optimization process. We thoroughly evaluate our methods on synthetic and real-world datasets, demonstrating 90+in depth and 20+color rendering performance and artifact-free novel view synthesis in long-term dynamic evolving scenes. The code and data are released as open-source.
@@
Appendix A Appendix
We provide implementation details, additional ablation studies, and qualitative results. This supplementary material includes rendered videos of reconstructed environments by GaME, compared to the baselines, on the Flat dataset, highlighting the effectiveness of our mapping approach. Additionally, we include implementation details, the ablation on sensitivity to noisy poses, and runtime analysis.
Implementation Details. In Tab. 7 we provide the hyperparameters used in our experiments.
Dataset $\lambda_{\mathrm{color}}$ $\lambda_{\mathrm{depth}}$ $\theta_{\mathrm{depth}}$ $\theta_{\mathrm{opacity}}$ $\theta_{\mathrm{color}}$ $\theta_{\mathrm{mask}}$ $\epsilon_{\mathrm{depth}}$ $\theta_{\mathrm{rotation}}$ $\theta_{\mathrm{translation}}$ $\theta_{\mathrm{ignore}}$
Flat [26] 1.0 1.0 4.0 0.3 0.1 0.4 2.0 50 5.0 0.9
Aria [20] 1.0 1.0 4.0 0.3 0.3 0.4 4.0 25 0.5 0.2
Table 7: Hyperparameters for Flat [26] and Aria [20] datasets.
GaME can handle noisy poses. We evaluate the camera poses on the Aria dataset with an external SLAM [18] system and use the estimated poses and keyframes instead of ground truth in Tab. 8. GaME is robust even when the pose estimation is not precise, exhibiting only a marginal performance drop under noisy pose conditions.
Camera Poses PSNR [dB] $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Depth L1 [cm] $\downarrow$ Ground-truth 26.63 / 26.71 0.920 / 0.920 0.179 / 0.180 2.4 / 2.4 Estimated 24.83 / 24.93 0.873 / 0.876 0.234 / 0.232 3.6 / 3.6
Table 8: Usage of noisy poses on Aria room0. GaME is robust to camera pose noise, exhibiting only a slight performance drop when using noisy poses instead of ground-truth. Camera poses and keyframes are estimated with an off-the-shelf reconstruction system [18].
Run-time Analysis. We compare the mapping runtime of our method with other state-of-the-art Gaussian-based systems in Tab. 9. Scenes with changing geometry naturally slow the mapping process, as existing methods struggle with seeding new geometry and optimizing under conflicting observations. In contrast, GaME is designed to explicitly handle such dynamic scene adaptation while being comparably efficient to the state-of-the-art mapping systems.
Metric SplaTAM [11] MonoGS [18] DG-SLAM [33] GaME (Ours)
[…]
Table 9: Runtime Analysis on Flat dataset. Evolving scenes significantly slow down the mapping process. While GaME supports scene adaptation mechanisms, it performs comparably with state-of-the-art mapping systems. The frame per second is calculated by dividing the time spent on mapping by the total number of processed frames. All metrics are profiled using an NVIDIA RTX 3090 GPU.
Note on Baselines. Wild-GS [39] cannot be run with ground-truth poses, as it relies on an off-the-shelf depth estimator that may be inconsistent with them. At the same time, it fails to track the camera poses on the Aria and Flat datasets. For this reason, we were not able to compare our method with it.

### v2

@@ 4.3 Keyframe Management
Since jointly optimizing Gaussians using all frames from a video stream is computationally infeasible, we maintain a smaller set of keyframes $W_{k}$. Effective keyframe management (KM) aims to select non-redundant keyframes that observe the same region while spanning a wide baseline to enforce stronger multiview constraints. Following [keetha2024splatam, yugay2023gaussianslam], keyframes are selected every time a frame exceeds a translation $\theta_{\text{translation}}\in\mathbb{R}$ or a rotation $\theta_{\text{rotation}}\in\mathbb{R}$ threshold.
@@
Covisibility. Upon triggering Add or Remove operations by DSA, all covisible keyframes are retrieved by reprojecting 3D points from the current frame and selecting frames where these points are not occluded. This approach differs from conventional 3DGS methods [murai2024monogs, xu2024dgslam], which determine covisibility based on the number of Gaussians used to re-render the keyframes. We found projection to be more reliable than re-rendering in multi-room environments, as Gaussians from different rooms may appear visible due to alpha blending, leading to erroneous covisibility estimates.
@@
Keyframe Masking. In evolving scenes, excluding outdated visual information plays a critical role, as the scene may change over time, making previous observations detrimental for the 3DGS optimization process. However, discarding entire keyframes that observe stale geometry can result in losing useful information. While some parts of the scene may change, other regions often remain stable and can provide valuable signals for 3DGS optimization. Therefore, rather than discarding entire keyframes as stale, GaME selectively ignores only the stale areas within them.
To achieve this, before re-initializing $\mathcal{G}_{\text{add}}$ or removing $\mathcal{G}_{\text{remove}}$ and $\mathcal{G}_{\text{remove}}^{\text{KF}}$, our keyframe management system renders them onto covisible keyframes and assesses conflicts via photometric and geometric losses Eq. 4. High-error regions correspond to the areas of the map that were changed but are not reflected in the observations. To mitigate the noise that typically happens in 3DGS renderings [kerbl2023gaussian_splatting], our KM module takes densely seeded SAM [kirillov2023segany] intersecting with the conflicting regions to produce areas that are aligned with the objects in the image and marks them as ignored. During optimization, all the keyframes that have SAM masks fully covered by the erroneous area are ignored. For the rest, GaME optimizes the losses from  Eq. 4 using the masked version of losses. We visualize the keyframes with masked out stale areas on Fig. 4.
Refer to caption Refer to caption
Figure 4: Visualizations of the outdated keyframes. After detecting the changes, our keyframe management masks out the areas of covisible keyframes observing changed geometry.
@@
SplaTAM [keetha2024splatam] MonoGS [murai2024monogs] DG-SLAM [xu2024dgslam] GaME (Ours) Ground Truth
@@ 5 Experiments
Datasets. We test our method on the Flat [schmid2022panoptic] dataset, which consists of two RGB-D sequences captured in a synthetic environment with significant changes occurring between. We further evaluate on the Aria [pan2023ariadigitaltwinnew] dataset to assess performance on real-world data. We select two recordings from two rooms each that have undergone long-term changes. Finally, the TUM-RGBD [sturm12iros] dataset is used to assess performance on three different static scenes following the protocol of [murai2024monogs].
Evaluation Metrics. To assess rendering quality, we compute PSNR, SSIM [wang2004image] and LPIPS [zhang2018unreasonable]. Rendering metrics on all the datasets are evaluated by rendering full-resolution images along the ground-truth trajectory. We assess the depth error using the L1 norm in centimeters. In the tables, results are color-coded to indicate ranking: best, second-best, and third-best.
Baselines. We compare GaME with state-of-the-art 3DGS online reconstruction systems MonoGS [murai2024monogs] and SplaTAM [keetha2024splatam]. We compare to DG-SLAM [lu20253dgs], a recent dynamic 3DGS method, to assess the ability of dynamic 3DGS SLAM to handle evolving scenes. Wild-GS [zheng2025wildgs] cannot be run with ground-truth poses, as it relies on an off-the-shelf depth estimator that may be inconsistent with them. At the same time, it fails to track the camera poses on the Aria and Flat datasets. For this reason, we were not able to compare our method with it.
@@ 5.1 Evolving Scene Mapping Results
We run GaME on the Flat and Aria datasets to evaluate rendering quality in evolving scenes, shown in Tabs. 1 and 2. The Flat dataset is designed to test mapping under scene changes and includes more substantial long-term dynamics. We note that all losses are computed over complete images, where changed areas typically occupy a smaller part. Nonetheless, we observe notable differences in rendering performance on both datasets, where GaME is the only method that accurately adapts the reconstruction to even subtle scene changes without compromising rendering quality. This granularity is shown in qualitative results in Fig. 5. Beyond high rendering quality, GaME is able to accurately resolve changes also for challenging cases such as small cutlery on the table and flat paintings on the walls.
Methods PSNR [dB] $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Depth L1 [cm] $\downarrow$
@@
[…]
DG-SLAM [lu20253dgs] 13.72 / 13.70 0.59 / 0.60 0.74 / 0.74 73.76 / 73.90
[…]
Table 1: Rendering performance on the Flat dataset. Our method significantly outperforms other baselines thanks to dynamic scene adaptaion and keyframe management system. Metrics are reported for input / novel views.
@@
DG-SLAM [lu20253dgs] room0 15.78 / 15.63 0.58 / 0.58 0.76 / 0.76 67.60 / 69.05
room1 12.62 / 12.44 0.65 / 0.64 0.70 / 0.71 55.47 / 57.03
Avg 14.20 / 14.04 0.62 / 0.61 0.73 / 0.74 61.54 / 63.04
@@
Table 2: Rendering performance on the Aria dataset. Our method significantly outperforms other baselines thanks to dynamic scene adaptaion and keyframe management system. Metrics are reported for input / novel views.
@@ 5.2 Ablation Studies
Add and Remove operations are important for Dynamic Scene Adaptation. We ablate the components of DSA in Tab. 3, showing that its presence significantly improves performance. While differences between individual DSA variants are less pronounced in final rendering quality, due to keyframe optimization correcting finer detail, combining both operations consistently yields the best performance.
Method PSNR [dB] $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Depth L1 [cm] $\downarrow$
No DSA 21.25 / 21.08 0.87 / 0.87 0.2 / 0.2 29.20 / 27.20
Add 23.34 / 23.07 0.92 / 0.91 0.15 / 0.15 11.00 / 11.20
Remove 24.13 / 23.89 0.93 / 0.93 0.14 / 0.14 9.50 / 8.20
Add and Remove (Ours) 24.55 / 24.26 0.93 / 0.93 0.14 / 0.14 6.9 / 7.9
Table 3: Dynamic Scene Adaptation (DSA) ablation on the Flat dataset. Combining Add and Remove operations gives the best performance. Metrics are reported for input / novel views.
@@
Method PSNR [dB] $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Depth L1 [cm] $\downarrow$
No KF Filtering 23.57 / 23.31 0.92 / 0.92 0.15 / 0.14 6.70 / 7.20
Full KF Filtering 13.72 / 13.34 0.58 / 0.57 0.56 / 0.55 764.40 / 763.30
Partial KF Filtering (Ours) 24.55 / 24.26 0.93 / 0.93 0.14 / 0.14 6.9 / 7.9
Table 4: Keyframe management ablation on the Flat dataset. Retaining keyframes is essential to constrain the background, but conflicting regions must be accurately filtered to avoid artifacts.
GaME can handle noisy poses. We evaluate the camera poses on the Aria dataset with an external SLAM [murai2024monogs] system and use the estimated poses and keyframes instead of ground truth in Tab. 5. GaME is robust even when the pose estimation is not precise, exhibiting only a marginal performance drop under noisy pose conditions.
Camera Poses PSNR [dB] $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Depth L1 [cm] $\downarrow$
Ground-truth 31.54 / 31.48 0.95 / 0.95 0.14 / 0.14 0.69 / 0.69
Estimated 31.45 / 31.53 0.95 / 0.95 0.14 / 0.14 0.70 / 0.70
Table 5: Effect of noisy poses on Aria room0. GaME is robust to camera pose noise, showing no significant impact on performance when using estimated poses instead of ground truth. Camera poses and keyframes are estimated with an off-the-shelf method [murai2024monogs].
GaME can reconstruct static scenes. While our method is specifically designed to reconstruct evolving scenes, it should not lose the ability to reconstruct static scenes. Results on the TUM-RGBD dataset shown in Tab. 6 show that GaME performs on par with state-of-the-art systems. This suggests that the introduced DSA and keyframe management are robust to noise and false positives in model updates and masking, which could deteriorate performance. In the future, GaME could well be combined with tracking and registration methods to improve performance, however, this is currently beyond the scope of this work.
@@
[…]
GaME is an online mapping system. While addressing scene changes requires additional computation time, our method performs on par with standard 3DGS reconstruction systems as shown in Tab. 7. While the run time is not yet real-time, our approach is of an incremental nature with partial optimization performed on every new keyframe. We believe that future improvements in 3DGS optimization speed will translate well to GaME without incurring notable extra cost to handle evolving scenes. Optionally, our approach lends itself to further final refinement. Both options are shown in Tab. 8. While refinement can potentially improve the final rendering quality, we note that the model is already well converged after incremental processing.
Metric SplaTAM [keetha2024splatam] MonoGS [murai2024monogs] DG-SLAM [xu2024dgslam] GaME (Ours)
[…]
Table 7: Runtime analysis on the Flat dataset. FPS is calculated by dividing the mapping time by the total number of frames. All metrics are profiled on an NVIDIA RTX 3090 GPU.
Method PSNR [dB] $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Depth L1 [cm] $\downarrow$
No Refinement 23.70 / 23.46 0.92 / 0.92 0.16 / 0.16 7.2 / 6.6
With Refinement 24.55 / 24.26 0.93 / 0.93 0.14 / 0.14 7.9 / 6.6
Table 8: Final refinement ablation on the Flat dataset. While refinement can further improve rendering quality, GaME already converges well during incremental processing.
GaME is robust. Since the main focus of our work is on the architectural challenges when handling evolving scenes in 3DGS and not image-based change detection, we opted for a comparably simple threshold-based approach. While GaME is readily extendable with more complex methods such as [lu20253dgs], we already find good performance and that GaME is robust w.r.t. critical hyperparameters. First, most of our hyperparameters have an interpretable meaning, making them intuitive and requiring minimal effort to adjust. For instance, to determine $\epsilon_{\text{depth}},\epsilon_{\text{color}}$, we inspected the rendering error in color and depth caused by adding a small object (e.g., a cup on the table) to the scene. For keyframe selection, we adopted standard rotation and translation thresholds [keetha2024splatam, yugay2023gaussianslam]. We experimentally show GaME to be robust w.r.t. hyperparameters by varying key parameters $\epsilon_{\text{depth}},\epsilon_{\text{color}},\epsilon_{\text{opacity}}$ with steps from the set of $\{-20\%,-10\%,10\%,20\%\}$ and report the mean and standard deviation of PSNR and Depth L1 error on the Flat dataset in Tab. 9. The minimal variation in scores confirms that our method does not rely on sensitive parameter tuning.
Hyperparameter PSNR PSNR Depth L1 [cm] Depth L1 [cm]
Mean (↑) Std. Dev. (↓) Mean (↓) Std. Dev. (↓)
$\epsilon_{\text{depth}}$ 31.52 0.21 0.68 0.03
$\epsilon_{\text{color}}$ 31.58 0.14 0.69 0.01
$\epsilon_{\text{opacity}}$ 31.56 0.11 0.68 0.01
Table 9: Hyperparameters robustness on the Aria room0 dataset. GaME is robust to hyperparameter variation.
@@ 6 Limitations
While GaME currently demonstrates robust rendering performance in evolving long-term dynamic scenes, several notable limitations exist. First, it does not yet handle short-term dynamic objects, which would result in inefficient addition and removal. Extending the dynamic scene adaptation mechanism to consider both dynamics is an exciting future direction. Second, we primarily study mapping with external pose tracking. Although sparse odometry systems can reject many changes as outliers [campos2021orb], change-aware tracking mechanisms are an exciting future direction. Finally, while GaME builds a consistent model of the scene at the final time, it would be interesting to explore a complete 4D representation of the history of all changes.
@@ 7 Conclusions
We presented GaME, the first online mapping system for evolving scenes with novel view synthesis capabilities. GaME utilizes novel dynamic scene adaptation operations to detect and correct conflicts in the incrementally built 3DGS model. Our keyframe management method furthermore appropriately ignores stale areas in keyframes, retaining useful information while correcting for changed regions, resulting in a well-conditioned 3DGS optimization process. We thoroughly evaluate GaME on synthetic and real-world datasets, demonstrating 29.7% improvement in PSNR, over $3\times$-improvement in L1 depth error, and artifact-free novel view synthesis in evolving scenes. The code and data are released open-source.